OpenStudy (anonymous):

constrained optimization involving multivariate functions A furniture company makes personal computer stands. For a certain model, the total cost (measured in thousands of dollars) when q hundred stands are produced is given by C(q) = 2q3 - 9q2 + 12q + 20 The company is currently capable of manufacturing between 75 and 600 (inclusive) per week. Determine the number of stands that must be produced so as to minimise total cost

OpenStudy (radar):

dC/dq=6q^2-18q+12 set 6q^2-18q+12+0 divide thru by 6 q^2-3q+2=0 solve by factoring (q-2)(q-1)=0 solve for q q=2 q=1 test each solution for minimum 2q-3=0 taking derivative of the derivative 2(2) -3 = 1 is positive, so is a minimum answer is 200 units

OpenStudy (anonymous):

were suppose to use Lagrange multiplier and probably Kuhn tucker conditions but thank you

OpenStudy (radar):

O.K. I gotta run, eye exam! Good luck.